OpenStudy (anonymous):

The midpoint of a line segment is located at (3, -2). If one of the endpoints is (1, 6), what is the other endpoint? Express your answer as an ordered pair.

OpenStudy (anonymous):

(5,-10)

OpenStudy (anonymous):

us midpoint formula

OpenStudy (anonymous):

You got the mid point (3,-2) And the mid point formula is. (x₁+x₂)/2,(y₁+y₂)/2 And setting one of the end point that is (1,6), and setting equal to the mid point would be something like this. (1+x₂)/2,(6+y₂)/2=(3,-2) So if you sett equal y equation, and x ecuation will be (1+x₂)/2=3 Solve for x x=5 And solving for y (6+y₂)/2=-2 y₂=-10 So the points are (5,-10) Hope you understand.